OpenStudy (anonymous):

Help!! Simplify 9√7-12√7

OpenStudy (anonymous):

what is 4 pineapples - 9 pineapples?

OpenStudy (anonymous):

-5?

OpenStudy (anonymous):

-5 what?

OpenStudy (anonymous):

@jagr2713 can you help with another?

jagr2713 (jagr2713):

\[Subtract 12\sqrt{7} from 9\sqrt{7} \to get -3\sqrt{7}\]

OpenStudy (anonymous):

6√10+9√10

jagr2713 (jagr2713):

what was the answer for the first one

OpenStudy (anonymous):

-3√7

jagr2713 (jagr2713):

so i got it right :D

OpenStudy (anonymous):

Yes! Thank you. Can you help me with the other one?

OpenStudy (anonymous):

@jagr2713 6√10+9√10